OpenStudy (anonymous):

complete factoring 2x^3+14x^2+12x

hero (hero):

2x(x^2 + 7x + 6) = 2x(x+6)(x+1) Give me another medal, lol

OpenStudy (anonymous):

-6x^6-12x^5+288x^4

hero (hero):

I gave you the answer to that already

hero (hero):

-6x^4(x^2+2x-48) = -6x^4(x+8)(x-6)

OpenStudy (anonymous):

a^8+5a^7b-6a^6b^2

OpenStudy (anonymous):

m^3n-10m^2n^2+24mn^3

hero (hero):

mn(m^2 - 10mn +24n^2) = mn(m - 6n)(m-4n)

hero (hero):

Okay, I'm going to bed

OpenStudy (anonymous):

wait one more!

OpenStudy (anonymous):

6x^3-4x^2+2x

hero (hero):

What the? You tricked me!

OpenStudy (anonymous):

last one i promise

OpenStudy (anonymous):

please help me with this last one

hero (hero):

This one appears to be complex

OpenStudy (anonymous):

whats tht mean

hero (hero):

That means it doesn't have any real roots

OpenStudy (anonymous):

so what did you get

OpenStudy (anonymous):

its a factor

hero (hero):

Well, I take that back....Here's the solution: \[2x(\sqrt{3}x-1)(\sqrt{3}x+1)\]

OpenStudy (anonymous):

thanks goodnight